Parth (parthkohli):

I want to learn about Sigma Notation. I know that it's used for finding the sum of a sequence. How do you write it?

OpenStudy (anonymous):

\[\text{\sum}\longrightarrow \sum\]

Parth (parthkohli):

No, not latex.. I mean, how exactly is it written?

Parth (parthkohli):

What is the number above the term and below?

OpenStudy (anonymous):

\[\large \sum_{n=1}^n n = 1+ 2+3+4+\ldots+n\]

Parth (parthkohli):

What is meant by this?

OpenStudy (anonymous):

It's sum of first \(n\) natural numbers.

OpenStudy (anonymous):

Rather, representation of sum of first \(n\) natural numbers.

Parth (parthkohli):

I did understand. But why is n equal to 1 + 2 + 3 + 4...n

Parth (parthkohli):

Shouldn't it be: \(\Large \color{purple}{\rightarrow \sum_{n = 1}^{n}(1 + 2 + 3 + 4....n) }\)

OpenStudy (anonymous):

Hmm \[\large 1 + \sum_{n=2}^n n= 1 + 2+\ldots +n\implies 1 + 2+3+\sum_{n=4}^n=1+2+\ldots +n \] Do you get it now?

OpenStudy (anonymous):

Sorry.

OpenStudy (anonymous):

It's supposed to be \[1+2+3+\sum_{n=4}^nn=1+2+\ldots+n\]

Parth (parthkohli):

Oops, that's even more confusing @Ishaan94

OpenStudy (anonymous):

If it were for your representation then it would be, \[\large\sum_{n=1}^n\left(1+2+3+\ldots+n\right) = 1 +\left(1+2\right)+\left(1+2+3\right)+\ldots +\left(1+2+\ldots +n\right)\]

Parth (parthkohli):

I mean what do you mean by the n = 4 in the bottom, and n on the top?

OpenStudy (anonymous):

argh, bad latex :/

Parth (parthkohli):

lol :D

OpenStudy (anonymous):

Hmm the \(n\) is evaluate from 4 to \(n\). 4 is the initial value and \(n\) is the final value.

OpenStudy (anonymous):

I am a bad teacher :/

Parth (parthkohli):

Oh, I'm getting it.

Parth (parthkohli):

So, it's the sum of values when you substitute n with 4, then 5, then 6, then 7 till n.

OpenStudy (anonymous):

Let f(n) be any function or expression.\[\sum_{n=1}^nf(n)\]Then,\[\sum_{n=1}^n f(n)=f(1) + f(2)+\ldots +f(n)\]

Parth (parthkohli):

Don't say that. I'm a bad learner.

Parth (parthkohli):

I am getting it... wow, that makes sense!

OpenStudy (anonymous):

"So, it's the sum of values when you substitute n with 4, then 5, then 6, then 7 till n." Yes. Exactly.

OpenStudy (lgbasallote):

it's been discussed to me countless times i still dont get it =_= i never boother nowadays

Parth (parthkohli):

So, wait, let me try to get the sum of first n natural numbers.

OpenStudy (lgbasallote):

learn about mathematical induction @ParthKohli it's fun :P

OpenStudy (anonymous):

For example, \[\sum_{n=1}^nn^2 =1 + 2^2 + 3^2 + \ldots + n^2 \]

OpenStudy (anonymous):

Yes, you should learn Mathematical Induction. It's fun. and at times challenging.

Parth (parthkohli):

 Is this the sum of first 4 natural numbers?